Last week a company's income was $530. from this week, $50 was withheld for expenses. the rest was shared by two partners in the ratio 3 to 5. how much did each receive?
Answer by checkley71(8403)

(530-50)=480 is the amount to divide in the 3:5 ratio.
3:5 is an share system
480/8=60 per share
3*60=180 for the smaller share.
5*60=300 for the larger share.
proof
180+300=480 the amount to be divided.